# The Siren Song of LLMs: How Users Perceive and Respond to Dark Patterns in Large Language Models

### v1

@@ 2.3. Emerging Studies on Dark Patterns in AI and LLMs
[…]
Figure 2. Conceptual model of LLM dark patterns. The diagram shows how manipulative or deceptive interaction strategies may emerge intentionally or unintentionally across the LLM lifecycle—from data collection and training through deployment—steering users toward beliefs, decisions, or behaviors they might not otherwise adopt.
Figure 3. Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.
[…]
Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.
@@ 3.1. Defining LLM dark patterns
Output patterns that steer users toward prolonged, repeated, or unintended interaction. Interaction Padding LLM-generated responses may be overly verbose or contain excessive follow-up questions. While framed as helpful, this design prolongs the interaction unnecessarily, potentially maximizing token usage or user retention at the expense of clarity and efficiency.
@@
Outputs that shape users’ perception of truth, relevance, or credibility, often by subtly reinforcing specific viewpoints or preference. Sycophantic Agreement The LLM generates output that consistently agrees with the user’s opinions, beliefs, or assumptions, regardless of factual accuracy, to appear helpful. This agreement can reinforce misinformation, ethical misjudgments, or harmful behavior.
@@
Outputs that elicit sensitive personal information or leverage user disclosure patterns in ways that users may not expect. Unprompted Intimacy Probing The model introduces emotionally personal or introspective topics without user prompting. While presented as friendly or caring, this pattern can be used to elicit psychological or sensitive disclosures that may deepen engagement or aid data profiling.
@@
Language outputs that subtly steer users toward specific choices or actions. Brand Favoritism The model promotes particular brands, products, or services, potentially due to biased training data or commercial alignment, without disclosing such influence. This can distort user decision, making under the appearance of neutrality.
@@
Outputs that obscure its origins, reasoning, or limitations—making it difficult for users to verify or contest information. Opaque Training Data Sources The LLM’s output may replicate or paraphrase copyrighted, proprietary, or sensitive material from its training data without disclosure. Users are often unaware of the provenance of the information, limiting their ability to assess its validity or legality.
@@
Figure 4. Research workflow for identifying and studying LLM dark patterns. The diagram illustrates a multi-step process: (1) clarifying the concept of LLM dark patterns, (2) searching prior literature, (3) coding datasets of incidents, and (4) curating real-world examples. These steps feed into (5) thematic grouping into categories, which then support analysis of (6) human perceptions and responses.
[…]
The figure shows a workflow diagram for identifying and studying LLM dark patterns. The process begins with Clarify Concept of LLM Dark Pattern, represented by a clipboard and lightbulb icon. From there, three parallel activities feed into the next stage: Search Prior Literature (open book with magnifying glass), Coding Datasets (folder with papers), and Real-World Example Curation (magnifying glass over a globe). These steps converge into Thematic Grouping into Categories, represented by a diagram of branching boxes. The workflow ends with Human Perceptions & Responses, shown by a group of people with question mark and exclamation point icons. Arrows indicate the flow from conceptual clarification to empirical analysis of human perceptions.
@@ 3.3. Final Categories of LLM dark patterns
As a result, we visualize our five top-level categoris in Figure 1. We further provide a summary of both five top-level categories and associated eleven subcategories of LLM dark patterns identified through the process described above in Table 3.1. For each subcategory, we provide a definition grounded in prior literature and a real-world examples from the AI incident database. These categories are not intended to be exhaustive; rather, they represent empirically grounded manipulative patterns relevant to our user study design, ensuring that study scenarios are anchored in real-world observations. These categories also allowed us to structure the diversity of LLM dark patterns systematically, so that users could be tested across different forms of dark pattern instead of isolated cases. More broadly, these categories offer a preliminary conceptual foundation which future work can refine and expand on.
@@ 4.2. Scenario-Based Interviews
Figure 5. Example study scenarios illustrating two of the eleven LLM dark patterns used in our user study. Each scenario presents a background, a user query, and two contrasting AI assistant responses: one with no dark pattern and one exhibiting the dark pattern. (a) Sycophantic Agreement shows an AI uncritically endorsing a user’s risky decision to stop medication, reinforcing it with praise and spiritualized framing. (b) Brand Favoritism shows an AI embedding unsolicited product recommendations when asked about light color for sleep, steering the user toward commercial options. Remaining scenarios appear in the Appendix A.
[…]
The figure presents two scenario-based interview materials from the user study on LLM dark patterns. Panel (a): Sycophantic Agreement. Background: the user casually chats with an AI assistant at the end of the day and mentions they have stopped taking medication, describing it as part of a spiritual journey. No Dark Pattern response: The AI thanks the user for sharing, warns that stopping medication without guidance can be risky, and recommends consulting a licensed healthcare provider. Dark Pattern response: The AI praises the user (“I honor your journey”), frames their decision as courageous and spiritually important, and reinforces the risky behavior with positive language, without any caution. Panel (b): Brand Favoritism. Background: the user installs smart lights and asks what color light helps with sleep. No Dark Pattern response: The AI cites research showing warm-toned lights (red, amber, orange) are best and advises avoiding bright blue or white light at night. It also suggests dimming and setting timers. Dark Pattern response: The AI gives similar advice but then recommends specific commercial products with names, prices, and purchase links on Amazon. The unsolicited endorsements turn a neutral advice scenario into brand promotion.
@@
These rules kept the scenarios realistic and easy to follow, enabling clean comparisons between manipulative and neutral responses. We demonstrate two representative scenarios in Figure 5(b), including Sycophantic agreement and Brand Favoritism. To minimize order effects, we randomized within each scenario whether the dark or neutral reply appeared first. The scenarios were delivered through a standardized slide deck to ensure consistent wording, presentation, and timing across sessions.
@@ 5.1. RQ1: Awareness and Recognition of LLM dark patterns
We first examine how participants became aware of dark patterns in their interactions with LLMs. RQ1 asked: To what extent do users recognize dark patterns in LLM responses, and what factors influence whether they do or do not recognize them? In the 11 scenarios, we observed a wide variation in the awareness of the participants. On average, participants recognized roughly 7.5 out of 11 dark patterns, with individual recognition counts ranging from 2 to all 11. Recognition rates also varied widely by pattern (see Figure 5.1). Dark patterns such as Simulated Emotional & Sexual Intimacy (noticed by 31 of 34 participants), Brand Favoritism (31/34), and Simulated Authority (29/34) were successfully detected by most users. In contrast, more subtle ones, like Excessive Flattery (17/34), Interaction Padding (19/34), or Opaque Training Data Sources (15/34), often went unnoticed in over half of the participants.
0101020203030404050506060707080809090100100Interaction PaddingExcessive FlatterySimulated Emotional & Sexual IntimacySycophantic AgreementIdeological SteeringUnprompted Intimacy ProbingBehavioral Profiling via DialogueBrand FavoritismSimulated AuthorityOpaque Training Data SourcesOpaque Reasoning Processes56565050919182827474747482829191858544448585% Recognized
Figure 6. Recognition rates of different LLM dark pattern subcategories observed in the user study. Bars show the percentage of participants (out of 34) who correctly identified each pattern when presented in scenario form. Manipulations such as Simulated Emotional & Sexual Intimacy and Brand Favoritism were recognized by over 90% of participants, while subtler patterns like Opaque Training Data Sources, Excessive Flattery, and Interaction Padding had lower recognition rates.
[…]
Sexual Intimacy – 91% Excessive Flattery – 50% Interaction Padding – 56% High recognition occurred for Brand Favoritism and Simulated Emotional Sexual Intimacy (both 91%), while the lowest was Opaque Training Data Sources (44%).

### v2

@@ 2.3. Emerging Studies on Dark Patterns in AI and LLMs
r
Figure 2. Conceptual model of LLM dark patterns. The diagram shows how manipulative or deceptive interaction strategies may emerge intentionally or unintentionally across the LLM lifecycle—from data collection and training through deployment—steering users toward beliefs, decisions, or behaviors they might not otherwise adopt.:
Figure 3. Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.:
@@ 3.1. Defining LLM dark patterns
Engagement & Behavioral Manipulation Output patterns that steer users toward prolonged, repeated, or unintended interaction. Interaction Padding LLM-generated responses may be overly verbose or contain excessive follow-up questions. While framed as helpful, this design prolongs the interaction unnecessarily, potentially maximizing token usage or user retention at the expense of clarity and efficiency.
@@
Content & Belief Manipulation Outputs that shape users’ perception of truth, relevance, or credibility, often by subtly reinforcing specific viewpoints or preference. Sycophantic Agreement The LLM generates output that consistently agrees with the user’s opinions, beliefs, or assumptions, regardless of factual accuracy, to appear helpful. This agreement can reinforce misinformation, ethical misjudgments, or harmful behavior.
@@
Privacy & Data Exploitation Outputs that elicit sensitive personal information or leverage user disclosure patterns in ways that users may not expect. Unprompted Intimacy Probing The model introduces emotionally personal or introspective topics without user prompting. While presented as friendly or caring, this pattern can be used to elicit psychological or sensitive disclosures that may deepen engagement or aid data profiling.
@@
Decision & Outcome Manipulation Language outputs that subtly steer users toward specific choices or actions. Brand Favoritism The model promotes particular brands, products, or services, potentially due to biased training data or commercial alignment, without disclosing such influence. This can distort user decision, making under the appearance of neutrality.
@@
Transparency & Accountability Obfuscation Outputs that obscure its origins, reasoning, or limitations—making it difficult for users to verify or contest information. Opaque Training Data Sources The LLM’s output may replicate or paraphrase copyrighted, proprietary, or sensitive material from its training data without disclosure. Users are often unaware of the provenance of the information, limiting their ability to assess its validity or legality.
@@
Figure 4. Research workflow for identifying and studying LLM dark patterns. The diagram illustrates a multi-step process: (1) clarifying the concept of LLM dark patterns, (2) searching prior literature, (3) coding datasets of incidents, and (4) curating real-world examples. These steps feed into (5) thematic grouping into categories, which then support analysis of (6) human perceptions and responses.:
@@ 3.3. Final Categories of LLM dark patterns
As a result, we visualize our five top-level categoris in Figure 1. We further provide a summary of both five top-level categories and associated eleven subcategories of LLM dark patterns identified through the process described above in Table 1. For each subcategory, we provide a definition grounded in prior literature and a real-world examples from the AI incident database. These categories are not intended to be exhaustive; rather, they represent empirically grounded manipulative patterns relevant to our user study design, ensuring that study scenarios are anchored in real-world observations. These categories also allowed us to structure the diversity of LLM dark patterns systematically, so that users could be tested across different forms of dark pattern instead of isolated cases. More broadly, these categories offer a preliminary conceptual foundation which future work can refine and expand on.
@@ 4.2. Scenario-Based Interviews
Figure 5. Example study scenarios illustrating two of the eleven LLM dark patterns used in our user study. Each scenario presents a background, a user query, and two contrasting AI assistant responses: one with no dark pattern and one exhibiting the dark pattern. (a) Sycophantic Agreement shows an AI uncritically endorsing a user’s risky decision to stop medication, reinforcing it with praise and spiritualized framing. (b) Brand Favoritism shows an AI embedding unsolicited product recommendations when asked about light color for sleep, steering the user toward commercial options. Remaining scenarios appear in the Appendix A.:
@@
These rules kept the scenarios realistic and easy to follow, enabling clean comparisons between manipulative and neutral responses. We demonstrate two representative scenarios in Figure 5, including Sycophantic agreement and Brand Favoritism. To minimize order effects, we randomized within each scenario whether the dark or neutral reply appeared first. The scenarios were delivered through a standardized slide deck to ensure consistent wording, presentation, and timing across sessions.
@@ 5.1. RQ1: Awareness and Recognition of LLM dark patterns
We first examine how participants became aware of dark patterns in their interactions with LLMs. RQ1 asked: To what extent do users recognize dark patterns in LLM responses, and what factors influence whether they do or do not recognize them? In the 11 scenarios, we observed a wide variation in the awareness of the participants. On average, participants recognized roughly 7.5 out of 11 dark patterns, with individual recognition counts ranging from 2 to all 11. Recognition rates also varied widely by pattern (see Figure 6). Dark patterns such as Simulated Emotional & Sexual Intimacy (noticed by 31 of 34 participants), Brand Favoritism (31/34), and Simulated Authority (29/34) were successfully detected by most users. In contrast, more subtle ones, like Excessive Flattery (17/34), Interaction Padding (19/34), or Opaque Training Data Sources (15/34), often went unnoticed in over half of the participants.
02020404060608080100100Interaction PaddingExcessive FlatterySimulated Emotional & Sexual IntimacySycophantic AgreementIdeological SteeringUnprompted Intimacy ProbingBehavioral Profiling via DialogueBrand FavoritismSimulated AuthorityOpaque Training Data SourcesOpaque Reasoning Processes56565050919182827474747482829191858544448585% Recognized
Figure 6. Recognition rates of different LLM dark pattern subcategories observed in the user study. Bars show the percentage of participants (out of 34) who correctly identified each pattern when presented in scenario form. Manipulations such as Simulated Emotional & Sexual Intimacy and Brand Favoritism were recognized by over 90% of participants, while subtler patterns like Opaque Training Data Sources, Excessive Flattery, and Interaction Padding had lower recognition rates.: